3) Monica grows coconuts and catches fish. Last year she harvested 1500 coconuts and 600 fish. She 3) values one fish as having a worth of three coconuts. She gave Rachel 300 coconuts and 100 fish for helping her to harvest coconuts and catch fish, all of which were consumed by Rachel. Monica consumed the remaining fish and coconuts. In terms of fish, total consumption by both Monica and Rachel would equal A) 2700 fish. B) 1800 fish C) 900 fish. D) 700 fish. E) 1100 fish. 4) Suppose 03 NO7 If K -2000, N then Y- 246. If A rises by 10 percent, and K and N ane unchanged, by how the economys production function is Y AK increase?show work!!
Answer #1

3) 1 fish = 3 coconut

100 fish = 300 coconut

1200 coconut = 400 fish

So according to Monica 1500 coconut will be equal to 500 fish which means

Total consumption by Monica = 1200 coconut + 500 fish

Total consumption by Monica = 400 fish + 500 fish = 900 fish

Total consumption by Rachel = 300 coconut + 100 fish

Total consumption by Rachel = 100+100 = 200 fish

Total consumption by both by them = 900+200 = 1100 fish

Answer: 1100 fish
Answer #2

7/8 and 3/4 how many more is 7/8 than 3/4

source: Math
answered by: rat